# Doubly-Dynamic ISAC Precoding for Vehicular Networks: A Constrained Deep Reinforcement Learning (CDRL) Approach

### v1

@@ VI Conclusions
In this paper, we proposed a CDRL approach to tackle the ISAC precoding problem in scenarios where both communication and sensing exhibit time variation. By utilizing various state observations such as the historical positions and initial CSI, the proposed scheme effectively adapts to the double dynamics. Moreover, the enhancement of our training algorithm allows us to effectively address constrained problems and seamlessly adapt to varying user number. The proposed scheme showcases enhancements in performance boundaries within double dynamics, achieving a minimum of 2 dB performance gain in sensing while ensuring SE for communications. It also diminishes reliance on perfect prior information and computational complexity, making it a promising solution for supporting vehicular networks.

### v2

@@ VI Conclusions
In this paper, we proposed a CDRL approach to tackle the ISAC precoding problem in scenarios where both communication and sensing exhibit time variation. By utilizing various state observations such as the historical positions and initial CSI, the proposed scheme effectively adapts to the double dynamics. Moreover, the enhancement of our training algorithm allows us to effectively address constrained problems and seamlessly adapt to varying user numbers. The proposed scheme showcases enhancements in performance boundaries within double dynamics, achieving a minimum of 2 dB performance gain in sensing while ensuring SE for communications. It also diminishes reliance on perfect prior information and computational complexity, making it a promising solution for supporting vehicular networks.